OpenStudy (anonymous):

safir can choose how he gets paid for a job. He can be paid $1000 all at once, or he can earn $1 for the first day, $2 for the second, $4 for the third, and so on, so that each day's pay is double that of the previous day. What is the positive difference of the total amounts Samir can be paid for completing a 10 day job, based on these two compensation plans?

OpenStudy (anonymous):

That is a geometric sequence. Hence, Amount if he goes $1 for day 1 etc. Amount = (1)(2)^(10-1) = 512 Positive difference = 1000 - 512 = 488